# GM-MoE: Low-Light Enhancement with Gated-Mechanism Mixture-of-Experts

### v2

@@ 5 Conclusion
In this paper, we present GM-MoE, a novel framework for low-light image enhancement that tackles three major challenges: color distortion, detail loss, and insufficient illumination, while ensuring strong generalization across diverse data domains. To achieve this, GM-MoE dynamically balances the contributions of three specialized sub-networks based on input features. Additionally, we introduce the SFEB module to enhance feature extraction. Extensive quantitative and qualitative evaluations on five benchmarks demonstrate that GM-MoE surpasses existing methods in PSNR and SSIM. In future work, we aim to extend GM-MoE for real-time image enhancement and further enhance its adaptability for real-world applications.
@@ A.1 Visualization of the results of a comparison experiment
(h) Ours (GM-MOE)
Figure 9: Visual comparison on the Huawei dataset. The models compared include RetinexNet, DeepUPE, Restormer, SNR-Net, Retinexformer, LightenDiffusion, Ours (our model), and Ground Truth. Among them, GM-MOE achieves better enhancement compared to other models. Zoom in to see more details of the differences between models.
@@
Figure 11: Visual comparison on the Huawei and Nikon datasets. The enhancement effect of GM-MOE is closer to the original picture than other networks.
Methods LOL-v1 [37] LOL-v2-real [45] LOL-v2-Sy [45] LSRW-Huawei [63] LSRW-Nikon [63] Params (M)
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
3DLUT[66] 14.35 0.445 17.59 0.721 18.04 0.800 18.12 0.6588 17.81 0.6289 0.59
Sparse[45] 17.20 0.640 20.06 0.816 22.05 0.905 20.33 0.6985 20.19 0.6571 1.08
RUAS[67] 18.23 0.720 18.37 0.723 16.55 0.652 20.46 0.7044 20.88 0.6637 1.03
DRBN [68] 20.13 0.830 20.29 0.831 23.22 0.927 20.61 0.7101 21.07 0.6704 1.83
Table 4: Performance comparison of selected methods on different datasets.
@@
In the Tab. 4, we also compare several low-light enhancement networks with GM-MOE. GM-MOE achieved significant PSNR improvements on all five datasets, especially on the LOL-v1 [37] and LOL-v2-Sy [45] datasets. Specifically, on the LOL-v1 dataset, GM-MOE achieved a PSNR of 26.66 decibels and an SSIM of 0.857, which is 6.53 decibels and 0.027 higher, respectively, than the best-performing network in the table, DRBN. On the LOL-v2-Sy dataset, GM-MOE achieved a PSNR of 26.30 decibels and SSIM is 0.937, which is 3.08 dB and 0.010 higher than DRBN, respectively. These results, as well as the improved SSIM values for all datasets, indicate that GM-MOE is more adaptable when processing images from different data domains.
@@ A.2 About the dataset
A.3 Code Availability
To support academic research, we will make the code public after the paper is accepted. The source code will be hosted on GitHub, and the link to the shared code base will be shared after publication.
A.4 Limitation and future works

### v5

@@ 5 Conclusion
This paper proposes the GM-MoE framework, which dynamically balances three expert subnetworks based on input image features, simultaneously addressing color bias, detail loss, and insufficient illumination in low-light images while ensuring strong generalization across diverse data domains. Extensive quantitative and qualitative evaluations on five benchmark datasets demonstrate that GM-MoE outperforms existing methods in both PSNR and SSIM metrics. Future work will focus on real-time enhancement and adaptive optimization.
[…]
Supplementary Material
@@ A.1 Visualization of the results of a comparison experiment
(h) PairLIE [9]
(i) SCI [36]
(j) QuadPrior [49]
(k) NeRCo [59]
(l) Ours (GM-MOE)
Figure 9: Visual comparison on the LSRW-Huawei dataset [14]. The models compared include RetinexNet, DeepUPE, Restormer, SNR-Net, Retinexformer, LightenDiffusion, PairLIE, SCI, QuadPrior, NeRCo, Ours (our model), and Ground Truth. Among them, GM-MOE achieves better enhancement compared to other models. Zoom in to see more details of the differences between models.
@@
Figure 11: Visual comparison on the Nikon dataset. The enhancement effect of GM-MOE is closer to the original picture than other networks.
@@ A.2 About the dataset
A.3 Generalisation to SID Dataset
On the SID benchmark [4], GM–MoE attains 24.80 PSNR and 0.69 SSIM, demonstrating strong performance.
A.4 Comparison with Competitive Baselines on LOL‑v1
As shown in Tab. 4, our method achieves a PSNR of 26.66 dB, an SSIM of 0.86, and a LPIPS of 0.098.
Methods PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
SCI [36] 14.78 0.53 0.392
NeRCo [59] 22.95 0.79 0.311
DiffLLE [60] 22.24 0.79
LightenDiffusion [19] 20.45 0.80 0.192
Ours 26.66 0.86 0.098
Table 4: Quantitative comparison on the LOL-v1 dataset among different methods: SCI [36], NeRCo [59], DiffLLE [60] , LightenDiffusion [19], and Ours.
A.5 Perceptual Quality (LPIPS)
As shown in Tab. 5 , we report the LPIPS scores of different models on different benchmarks.
Method LOL‑v1 LOL‑v2‑Real LOL‑v2‑Synthetic
Retinexformer [2] 0.129 0.171 0.059
LLFormer [67] 0.167 0.211 0.066
GM–MoE (Ours) 0.098 0.100 0.041
Table 5: LPIPS ($\downarrow$) comparison across three LOL datasets. Lower is better.
A.6 How are expert modules coordinated?
As shown in Fig. 12 , we enforce the use of only one expert at a time. The heatmap of Expert1 exhibits yellow-to-red contributions in fringe details, indicating its primary role in local chroma restoration. Expert2’s heatmap shows predominantly dark blue regions with only faint highlights in limited detailed areas, suggesting its specialization in fine texture recovery. Expert3 demonstrates extensive orange-yellow coverage across both the main fringe and background regions. These three heatmaps reveal spatial complementarity among the experts, enabling the final enhanced results to achieve both rich detail preservation and more realistic color reproduction.
Figure 12: The heatmaps show distinct focus areas across the three experts, forming a coherent, synergistic attention distribution.
A.7 Computational Efficiency
Our method delivers superior low‑light enhancement quality while incurring a computational cost of 27.2 GFLOPs, which is still practical for real‑time deployment. Future work will focus on further reducing latency and memory footprint without sacrificing restoration accuracy.
A.8 Ablation study on network structure and expert interactions
To explore in depth the relationship between the number of parameters and performance, we conducted three sets of experiments: fixed weights, cascaded networks, and constrained parameter growth.
No. Variant Params (M) PSNR ($\uparrow$) SSIM ($\uparrow$)
1 Original GM-MoE 19.99 23.65 0.80
2 Without dynamic gating (fixed weights) 19.86 21.70 0.71
3 Serializing three expert networks 19.86 21.34 0.83
4 Experts’ channels concat + $1\!\times\!1$ fusion 20.60 17.84 0.70
Table 6: Ablation study on network structure and expert interactions evaluated on the LOLv2-Real dataset.
A.9 Limitation and future works
@@
Acknowledgments
This work was supported by the Xinjiang Uygur Autonomous Region Tianshan Ying Talents Leading Talents Program for Scientific and Technological Innovation (2023TSYCLJ0025), the National Natural Science Foundation of China (No. 62266044, 61563052, 61862061), and the National Key Research and Development Program of China (2021YFB2802100).